OpenStudy (anonymous):

What are the possible dimensions for a triangle with an area of 32 cm2? a. b = 4 cm h = 8 cm c. b = 8 cm h = 8 cm b. b = 2 cm h = 16 cm d. b = 6 cm h = 8 cm

OpenStudy (ivettef365):

A = 1/2 b.h

OpenStudy (ivettef365):

so multiply 32 x 2 = 64 now find which numbers will multiply to 64

OpenStudy (anonymous):

so the base and height would be 8?

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

now to check use the formula for are A = 1/2 bh A = 1/2 (8) (8) A = 1/2 (64) A = 32